Factor completely: a^4-1
I got the answer (a^2+1)(a+1)(a-1) but I don't know how
Please show me step by step to get the answer. thanks

Answer by jorel1380(3719)

a4-1=(a2+1)(a2-1)
""=(a2+1)(a+1)(a-1)..